OpenStudy (anonymous):

need help converting x^2+y^2=16 to polar form

OpenStudy (apples):

Note that \[r^2 = x^2 + y^2\] thus \[r^2 = 16\] Take the square root of both sides, and the answer becomes r = 4.

OpenStudy (anonymous):

alright so i got that, now about notating it into polar coordinates, would i just use the arctan formula?

OpenStudy (apples):

Hrm? It's already in polar form. You don't need to do anything else.

OpenStudy (anonymous):

ah, i c. well thank you. i will do a few examples with your assistance.